OpenStudy (anonymous):

give a substitution (not necessarily a trigonometric) which could be used to find the integral of x/sqrt{(x^2+10}

OpenStudy (anonymous):

try \(u=x^2+10\) and you will be done quickly

OpenStudy (anonymous):

,did you get what satellite is saying?

OpenStudy (anonymous):

yes thank you!

OpenStudy (anonymous):

integral of x/sqrt{(x^2+10} let u=(x^2+10) du=2xdx sub this integral of x/sqrt{(x^2+10} =(1/2)integral of 2xdx/sqrt{(x^2+10} =(1/2)integral of2u^(-1/2) du =(1/2)u^1/2 /(1/2) =sqrt(x^2+10) +C